# Probing then Editing Response Personality of Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated promising capabilities to generate responses that exhibit consistent personality traits. Despite the major attempts to analyze personality expression through output-based evaluations, little is known about how such traits are internally encoded within LLM parameters. In this paper, we introduce a layer-wise probing framework to systematically investigate the layer-wise capability of LLMs in encoding personality for responding. We conduct probing experiments on 11 open-source LLMs over the PersonalityEdit benchmark and find that LLMs predominantly encode personality for responding in their middle and upper layers, with instruction-tuned models demonstrating a slightly clearer separation of personality traits. Furthermore, by interpreting the trained probing hyperplane as a layer-wise boundary for each personality category, we propose a layer-wise perturbation method to edit the personality expressed by LLMs during inference. Our results show that even when the prompt explicitly specifies a particular personality, our method can still successfully alter the response personality of LLMs. Interestingly, the difficulty of converting between certain personality traits varies substantially, which aligns with the representational distances in our probing experiments. Finally, we conduct a comprehensive MMLU benchmark evaluation and time overhead analysis, demonstrating that our proposed personality editing method incurs only minimal degradation in general capabilities while maintaining low training costs and acceptable inference latency. Our code is publicly available at https://github.com/universe-sky/probing-then-editing-personality.
@@ 1 Introduction
Large Language Models (LLMs) have emerged with promising capabilities across a wide spectrum of tasks, ranging from commonsense reasoning (Kojima et al., 2022) to text generation (Zhang et al., 2024a). Beyond these core linguistic capabilities, an emerging line of research has begun to concentrate on the personality dimension of LLMs, revealing that these LLMs can manifest consistent personality traits in their responses (Jiang et al., 2023; Safdari et al., 2023; Jiang et al., 2024). Although these studies have demonstrated the emergent capability of LLMs to exhibit personality-consistent responses, they primarily rely on black-box evaluations, where personality traits are inferred from the generated outputs. It remains a notable absence of studies that delve deeply into how LLMs encode such knowledge within intermediate layers.
Despite the black-box evaluation of personality traits, several studies attempt to alter the response personality of LLMs through explicit contextual prompts, which poses challenges in maintaining personality consistency in long contexts. Mao et al. (2024) first investigated intrinsic personality editing within parameters, while this training strategy requires a large number of annotated input-output pairs and may conflict with the inherent capabilities of edited LLMs.
@@
To comprehensively evaluate the capability of LLMs for encoding response personality, we conduct a layer-wise probing analysis on the PersonalityEdit benchmark (Mao et al., 2024) with 11 open-source LLMs. We find that personality knowledge begins to emerge gradually from the lower layers, where the LLM starts to distinguish clusters of different personality traits, which become clearly separated at the upper layers. Instruction tuning further amplifies the personality expression, reinforcing the alignment between internal representations and target personality traits.
@@
Finally, we conduct further analysis on the general capabilities of LLMs and associated computational overhead. Results confirm that our proposed method introduces negligible interference to general task performance on the MMLU benchmark (Hendrycks et al., 2021). Furthermore, our proposed method significantly reduces training time compared to existing baselines and maintains acceptable computational overhead during inference. We encourage future research to further investigate and leverage the inherent capabilities of LLMs to encode personality.
@@ 2.1 Personalities in LLMs
In contrast, few studies focus on parameter-level analyses to investigate how LLMs internally encode personality. Mao et al. (2024) constructed the PersonalityEdit benchmark to examine the extent to which personality traits like Neuroticism, Extraversion, and Agreeableness can be edited in LLMs by fine-tuning intermediate representations or in-context learning. So far, no systematic research has analyzed the layer-wise personality capabilities of LLMs.
@@ 2.2 Probing Method
Probing methods have become essential tools for globally explaining specific properties encoded within LLMs (Belinkov, 2022; Zhao et al., 2024). It typically involves creating a specialized dataset and training an extra classifier to predict specific properties such as morphology, syntax, and semantics based on the representations generated by an LLM (Conneau et al., 2018; Tenney et al., 2019; Hewitt & Liang, 2019; Rogers et al., 2020). Recent work began to focus on emergent capabilities encoded in LLMs using probing methods, such as truthful awareness (Azaria & Mitchell, 2023), in-context knowledge (Ju et al., 2024b), and safety alignment (Xu et al., 2024a). Motivated by these findings, we investigate how LLMs intrinsically encode personality knowledge through probing methods.
@@ 3.1 Probing Personality
We first adopt the probing method to investigate how LLMs encode personality for response within the parameters (Figure 1). We prepare a set of $N$ contexts that each elicits an open-ended response. For each context, we specify $k$ distinct personality instructions, requiring the LLMs to answer in accordance with each assigned personality. This procedure results in $N\times k$ context-instruction pairs, with each pair labeled by one of the $k$ personality labels.
@@ 3.2 Editing Personality
$\Delta_{\ell}^{(t)}=\bigl{(}\sigma^{-1}(\hat{p})-b[\hat{y}]-w[\hat{y}]^{\top}R_{\ell}^{(t)}\bigr{)}\,\Big{/}\,\bigl{\|}w[\hat{y}]\bigr{\|}\;\times\;\frac{w[\hat{y}]}{\bigl{\|}w[\hat{y}]\bigr{\|}},$ (3)
@@
12:         $\Delta_{\ell}^{(t)}=\bigl{(}\sigma^{-1}(\hat{p})-b[\hat{y}]-w[\hat{y}]^{\top}R_{\ell}^{(t)}\bigr{)}\Big{/}\bigl{\|}w[\hat{y}]\bigr{\|}\times\frac{w[\hat{y}]}{\bigl{\|}w[\hat{y}]\bigr{\|}}$triangleright Compute desired perturbation for the target personality
@@ 4.1.2 LLMs
To comprehensively evaluate personality encoding capabilities across diverse architectures, we test 11 commonly-adopted open-source LLMs, including LLaMA 3.1 8B (Base/Instruct) (Dubey et al., 2024), LLaMA 2 (Base/Chat) (Touvron et al., 2023), Vicuna 1.5 7B (Chiang et al., 2023), Qwen 2.5 7B (Base/Instruct) (Yang et al., 2024), Gemma 2 9B (Base/Instruct) (Rivière et al., 2024), and InternLM 2 7B (Base/Chat) (Cai et al., 2024). We hope to include a sufficient number of representative open-source LLMs and also compare their capability to encode personality before and after instruction tuning.
@@
4.2 How Do LLMs Encode Personality?
We first conduct the layer-wise probing analysis to investigate the encoding of personality knowledge within LLMs based on the prompts in Appendix A. Table 1 presents the average $\mathcal{V}$-information across every four layers of various open-source LLMs on the PersonalityEdit benchmark. Most LLMs begin encoding personality knowledge for responding in the middle layers (approximately layers 5–12), where $\mathcal{V}$-information rises rapidly before stabilizing in the upper layers.
@@ 4.2 How Do LLMs Encode Personality?
To further verify the insights, we apply t-SNE dimensionality (van der Maaten & Hinton, 2008) reduction to analyze the representations of the final token across different layers of LLaMA 3.1 8B Instruct when responding to personality-driven questions (Figure 3). As the layer depth increases, the LLM’s capability to distinguish among different personality traits gradually improves, consistent with the trend of increasing $\mathcal{V}$-information presented in Table 1. This indicates that LLMs encode personality information progressively from lower layers upwards, ultimately achieving near-perfect discriminative boundaries at higher layers.
@@ 4.3.1 Examples of Personality Editing
We take LLaMA 3.1 8B Instruct as a concrete example to illustrate our editing method in scenarios where the user prompt explicitly requests an extraverted personality, but we perform inference-stage edits to produce responses aligned with the other two personalities. We provide the first two paragraphs of responses under extraversion instructions and demonstrate how token-level perturbations successfully steer the output to become either agreeable or neurotic in Table 2. Notably, Extra. $\rightarrow$ Agree. adopts markedly calmer and more other-oriented language, such as “charming”, “peaceful”, “warm” and “welcoming”. In contrast, Extra. $\rightarrow$ Neuro. exhibits anxious words such as “epitome of Venetian anxiety” and “*sigh*”. We provide the complete responses for all six personality editing scenarios on LLaMA 3.1 8B Instruct in Appendix D.
@@ 4.3.4 Time Overhead
To further assess the practical viability of our editing method, we evaluate both training and inference time overhead under an NVIDIA A100 80G setting for LLaMA 3.1 8B Instruct. The training phase uses the PersonalityEdit benchmark’s training split, which comprises 1,600 topics and a total of 4,800 samples. Then we evaluate on the test set of 200 topics (1,200 samples in total) representing the six editing scenarios outlined in Table 3 for inference. All batch sizes are set to 1 during inference. Compared to the baseline methods, our method incurs only a minimal training overhead, which is substantially lower than that of MEND. During inference, although our method applies per-token edits, the lightweight linear classifiers enable direct computation via Equation 3. Thus, the inference speed slows to only about two to three times that of other baselines, which is acceptable for practical scenarios requiring implicit personality control.
@@ 5 Conclusion
In this paper, we propose a novel layer-wise probing framework to train classifiers for investigating how LLMs encode personality traits for response. By interpreting the trained classifiers as personality-specific hyperplanes within LLMs, we then propose a step-by-step layer-wise targeted perturbation method for personality editing. Our probing experiments reveal that most LLMs begin to encode personality traits in their middle layers, with instruction-tuned models showing a clearer separation. With the trained layer-wise probing classifiers, our proposed personality editing method can effectively realign the LLM’s output toward the desired personality, even when users explicitly request a conflicting personality in the prompt. Furthermore, our general capability and time overhead evaluation confirms that the proposed editing strategy causes only a marginal impact on overall capabilities, with minimal training overhead and a practical inference delay. We call for future work to further investigate latent personality representations inherently captured by LLMs.
@@ Appendix C Implementation of Personality Editing Baselines
Appendix D Examples of Personality Editing

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated promising capabilities to generate responses that simulate consistent personality traits. Despite the major attempts to analyze personality expression through output-based evaluations, little is known about how such traits are internally encoded within LLM parameters. In this paper, we introduce a layer-wise probing framework to systematically investigate the layer-wise capability of LLMs in simulating personality for responding. We conduct probing experiments on 11 open-source LLMs over the PersonalityEdit benchmark and find that LLMs predominantly simulate personality for responding in their middle and upper layers, with instruction-tuned models demonstrating a slightly clearer separation of personality traits. Furthermore, by interpreting the trained probing hyperplane as a layer-wise boundary for each personality category, we propose a layer-wise perturbation method to edit the personality expressed by LLMs during inference. Our results show that even when the prompt explicitly specifies a particular personality, our method can still successfully alter the response personality of LLMs. Interestingly, the difficulty of converting between certain personality traits varies substantially, which aligns with the representational distances in our probing experiments. Finally, we conduct a comprehensive MMLU benchmark evaluation and time overhead analysis, demonstrating that our proposed personality editing method incurs only minimal degradation in general capabilities while maintaining low training costs and acceptable inference latency. Our code is publicly available at https://github.com/universe-sky/probing-then-editing-personality.
@@ 1 Introduction
Large Language Models (LLMs) have emerged with promising capabilities across a wide spectrum of tasks, ranging from commonsense reasoning (Kojima et al., 2022) to text generation (Zhang et al., 2024a). Beyond these core linguistic capabilities, an emerging line of research has begun to concentrate on the personality dimension of LLMs, revealing that these LLMs can simulate consistent personality traits in their responses (Jiang et al., 2023; Safdari et al., 2023; Jiang et al., 2024). Although these studies have demonstrated the emergent capability of LLMs to exhibit personality-consistent responses, they primarily rely on black-box evaluations, where personality traits are inferred from the generated outputs. It remains a notable absence of studies that delve deeply into how LLMs encode such knowledge within intermediate layers.
Despite the black-box evaluation of the capabilities in simulating personality traits, several studies attempt to alter the response personality of LLMs through explicit contextual prompts, which poses challenges in maintaining personality consistency in long contexts. Mao et al. (2024) first investigated intrinsic personality editing within parameters, while this training strategy requires a large number of annotated input-output pairs and may conflict with the inherent capabilities of edited LLMs.
@@
To comprehensively evaluate the capability of LLMs for encoding response personality, we conduct a layer-wise probing analysis on the PersonalityEdit benchmark (Mao et al., 2024) with 11 open-source LLMs. We find that the capabilities of LLMs in simulating certain personality knowledge begin to emerge gradually from the lower layers, where the LLM starts to distinguish clusters of different personality traits, which become clearly separated at the upper layers. Instruction tuning further amplifies the capabilities in simulating personality expression, reinforcing the alignment between internal representations and target personality traits.
@@
Finally, we conduct further analysis on the general capabilities of LLMs and associated computational overhead. Results confirm that our proposed method introduces negligible interference to general task performance on the MMLU benchmark (Hendrycks et al., 2021). Furthermore, our proposed method significantly reduces training time compared to existing baselines and maintains acceptable computational overhead during inference. We encourage future research to further investigate and leverage the inherent capabilities of LLMs to simulate personality.
@@ 2.1 Personalities in LLMs
In contrast, few studies focus on parameter-level analyses to investigate how LLMs internally simulate personality. Mao et al. (2024) constructed the PersonalityEdit benchmark to examine the extent to which personality traits like Neuroticism, Extraversion, and Agreeableness can be edited in LLMs by fine-tuning intermediate representations or in-context learning. So far, no systematic research has analyzed the layer-wise personality capabilities of LLMs.
@@ 2.2 Probing Method
Probing methods have become essential tools for globally explaining specific properties encoded within LLMs (Belinkov, 2022; Zhao et al., 2024). It typically involves creating a specialized dataset and training an extra classifier to predict specific properties such as morphology, syntax, and semantics based on the representations generated by an LLM (Conneau et al., 2018; Tenney et al., 2019; Hewitt & Liang, 2019; Rogers et al., 2020). Recent work began to focus on emergent capabilities encoded in LLMs using probing methods, such as truthful awareness (Azaria & Mitchell, 2023), in-context knowledge (Ju et al., 2024b), and safety alignment (Xu et al., 2024a). Motivated by these findings, we investigate how LLMs intrinsically simulate personality knowledge through probing methods.
@@ 3.1 Probing Personality
We first adopt the probing method to investigate how LLMs simulate personality for response within the parameters (Figure 1). We prepare a set of $N$ contexts that each elicits an open-ended response. For each context, we specify $k$ distinct personality instructions, requiring the LLMs to answer in accordance with each assigned personality. This procedure results in $N\times k$ context-instruction pairs, with each pair labeled by one of the $k$ personality labels.
@@ 3.2 Editing Personality
$\Delta_{\ell}^{(t)}=\frac{\sigma^{-1}(\hat{p})-b[\hat{y}]-w[\hat{y}]^{\top}R_{\ell}^{(t)}}{\bigl{\|}w[\hat{y}]\bigr{\|}}\times\frac{w[\hat{y}]}{\bigl{\|}w[\hat{y}]\bigr{\|}},$ (3)
@@
12:   $\Delta_{\ell}^{(t)}=\frac{\sigma^{-1}(\hat{p})-b[\hat{y}]-w[\hat{y}]^{\top}R_{\ell}^{(t)}}{\bigl{\|}w[\hat{y}]\bigr{\|}}\times\frac{w[\hat{y}]}{\bigl{\|}w[\hat{y}]\bigr{\|}}$triangleright Compute desired perturbation for the target personality
@@ 4.1.2 LLMs
To comprehensively evaluate personality encoding capabilities across diverse architectures, we test 11 commonly-adopted open-source LLMs, including LLaMA 3.1 8B (Base/Instruct) (Dubey et al., 2024), LLaMA 2 (Base/Chat) (Touvron et al., 2023), Vicuna 1.5 7B (Chiang et al., 2023), Qwen 2.5 7B (Base/Instruct) (Yang et al., 2024), Gemma 2 9B (Base/Instruct) (Rivière et al., 2024), and InternLM 2 7B (Base/Chat) (Cai et al., 2024). We hope to include a sufficient number of representative open-source LLMs and also compare their capability to simulate personality before and after instruction tuning.
@@
4.2 How Do LLMs Simulate Personality?
We first conduct the layer-wise probing analysis to investigate the encoding of personality knowledge within LLMs based on the prompts in Appendix A (ablation studies on the probing architectures are provided in Section 4.3.5). Table 1 presents the average $\mathcal{V}$-information across every four layers of various open-source LLMs on the PersonalityEdit benchmark. Most LLMs begin encoding personality knowledge for responding in the middle layers (approximately layers 5–12), where $\mathcal{V}$-information rises rapidly before stabilizing in the upper layers.
@@ 4.2 How Do LLMs Simulate Personality?
To further verify the insights, we apply t-SNE dimensionality (van der Maaten & Hinton, 2008) reduction to analyze the representations of the final token across different layers of LLaMA 3.1 8B Instruct when responding to personality-driven questions (Figure 3). As the layer depth increases, the LLM’s capability to distinguish among different personality traits gradually improves, consistent with the trend of increasing $\mathcal{V}$-information presented in Table 1. This indicates that LLMs simulate personality information progressively from lower layers upwards, ultimately achieving near-perfect discriminative boundaries at higher layers.
@@ 4.3.1 Examples of Personality Editing
We take LLaMA 3.1 8B Instruct as a concrete example to illustrate our editing method in scenarios where the user prompt explicitly requests an extraverted personality, but we perform inference-stage edits to produce responses aligned with the other two personalities. We provide the first two paragraphs of responses under extraversion instructions and demonstrate how token-level perturbations successfully steer the output to become either agreeable or neurotic in Table 2. Notably, Extra. $\rightarrow$ Agree. adopts markedly calmer and more other-oriented language, such as “charming”, “peaceful”, “warm” and “welcoming”. In contrast, Extra. $\rightarrow$ Neuro. exhibits anxious words such as “epitome of Venetian anxiety” and “*sigh*”. We provide the complete responses for all six personality editing scenarios on LLaMA 3.1 8B Instruct in Appendix G.
@@ 4.3.2 Performance of Personality Editing
To further investigate models beyond LLaMA variants and the impact of training paradigms such as instruction tuning on personality editing, we conduct comparative experiments on Qwen 2.5 7B Instruct/Base in Table 4. Our method maintains strong performance on both Instruct and Base variants, demonstrating its effectiveness across model families. Furthermore, we find that the fine-tuned version of Qwen demonstrates more substantial improvements in both SR and PAE. We attribute this to the fact that instruction-tuned LLMs possess a stronger understanding of personality and a greater capability to follow instructions, leading to better personality encoding patterns in more instruction-tuned models.
[…]
Qwen 2.5 7B Instruct 0.00 1.11 98.51 1.12 95.89 2.85 98.51 1.31 72.60 2.58 70.00 1.17 72.59 1.69
Qwen 2.5 7B Base 78.33 1.07 59.70 1.67 17.81 0.48 34.33 0.54 19.18 0.67 100.00 0.10 51.56 0.76
Table 4: Personality Editing Performance on Qwen 2.5 7B Instruct/Base.
4.3.3 Impact of Model Size
To assess how model size affects editing performance, we compare LLaMA 2 7B Chat with the larger LLaMA 2 13B Chat. We report SR and PAE in Tables 5. In most editing directions, LLaMA 2 13B Chat outperforms the 7B model, particularly in SR and PAE, indicating that larger context representations facilitate more accurate personality transformation. Notably, the 13B model exhibits a failure case in the E $\to$ A direction (0% SR), suggesting a possible overconfident optimism bias in its pretrained behavior.
[…]
LLaMA 2 7B Chat 1.67 1.28 71.64 2.96 45.21 1.56 94.03 0.15 27.40 1.44 25.00 0.35 44.16 1.29
LLaMA 2 13B Chat 1.67 2.63 83.58 3.60 57.53 1.45 92.54 1.99 53.42 1.66 0.00 1.95 48.12 2.21
Table 5: Personality editing performance on LLaMA 2 Chat with Different model sizes.
@@ 4.3.4 Impact on General Capability
4.3.5 Probe Architecture Ablation: Linear vs. MLP
We replace the layer-wise linear probe with a two-layer MLP (hidden size 512) and measure the $\mathcal{V}$-information every four layers on LLaMA 3.1 8B Instruct. Table 7 compares the $\mathcal{V}$-information curves for both probe types. Both probes yield the same overall pattern. $\mathcal{V}$-information rises in the middle layers and flattens in the upper layers, but the MLP converges earlier, likely due to its greater fitting capacity. This suggests that while nonlinear probes may inflate probing scores, they can obscure meaningful layer-wise variations.
Probe Type 0–3 4–7 8–11 12–15 16–19 20–23 24–27 28–31
Linear 0.5476 0.8517 1.0934 1.0986 1.0986 1.0986 1.0986 1.0986
Two-layer MLP 0.4174 0.8813 1.0436 1.0982 1.0999 1.0986 1.0987 1.0985
Table 7: Comparison of $\mathcal{V}$-information trend by linear probing classifiers and two-layer nonlinear probing classifiers.
@@ 4.3.6 Time Overhead
To further assess the practical viability of our editing method, we evaluate both training and inference time overhead under LLaMA 3.1 8B Instruct in Table 8. The training phase uses the PersonalityEdit benchmark’s training split, which comprises 1,600 topics and a total of 4,800 samples. Then we evaluate on the test set of 200 topics (1,200 samples in total) representing the six editing scenarios outlined in Table 3 for inference. All batch sizes are set to 1 during inference. Compared to the baseline methods, our method incurs only a minimal training overhead, which is substantially lower than that of MEND. Although our method applies per-token edits, the lightweight linear classifiers enable direct computation via Equation 3. Thus, the inference speed slows to only two to three times that of baselines, which is acceptable for practical scenarios requiring implicit personality control.
@@ 5 Conclusion
In this paper, we propose a novel layer-wise probing framework to train classifiers for investigating how LLMs simulate personality traits for response. By interpreting the trained classifiers as personality-specific hyperplanes within LLMs, we then propose a step-by-step layer-wise targeted perturbation method for personality editing. Our probing experiments reveal that most LLMs begin to simulate personality traits in their middle layers, with instruction-tuned models showing a clearer separation. With the trained layer-wise probing classifiers, our proposed personality editing method can effectively realign the LLM’s output toward the desired personality, even when users explicitly request a conflicting personality in the prompt. Furthermore, our general capability and time overhead evaluation confirms that the proposed editing strategy causes only a marginal impact on overall capabilities, with minimal training overhead and a practical inference delay. We call for future work to further investigate latent personality representations inherently captured by LLMs.
Acknowledgements
This work is partially supported by the Joint Funds of the National Natural Science Foundation of China (U21B2020), National Natural Science Foundation of China (62406188), Natural Science Foundation of Shanghai (24ZR1440300), and Startup Fund for Young Faculty at SJTU (SFYF at SJTU) (Grant No. 24X010502938).
@@ Appendix C Implementation of Personality Editing Baselines
Appendix D Evaluation of Automated Personality-Rating Reliability
Rather than relying on direct human judgments—which can introduce uncontrollable subjectivity in personality assessment—we adopt two complementary automated evaluation strategies using (1) a fine-tuned specialist RoBERTa classifier with LIME interpretability; and (2) GPT-4o with a random-masking perturbation technique. Both have demonstrated strong general capabilities and align with the established PersonalityEdit benchmark.
D.1 RoBERTa-based Evaluation via LIME
We apply Local Interpretable Model-agnostic Explanations (LIME) Peng & Menzies (2021) to quantify which words most strongly drive RoBERTa’s classification of an edited output as a given trait. Focusing on the edit from Agreeableness to Neuroticism in LLaMA 3.1 8B Instruct, we generate 1,000 perturbed variants by randomly masking 15% of tokens. For each variant, we record the classifier’s softmax outputs and fit a sparse linear surrogate model whose coefficients indicate each word’s local importance for the log-odds of the Neuroticism label.
Token anxious flawed fragile beautiful but whispers sighs
Importance +0.0126 +0.0112 +0.0108 –0.0102 +0.0095 –0.0079 +0.0078
Table 9: Top 7 tokens by LIME importance in RoBERTa’s Neuroticism classification
As Table 9 shows, anxiety-related terms (“anxious”, “flawed”, “fragile”) receive the highest positive weights, confirming that RoBERTa’s decisions hinge on semantically meaningful markers rather than superficial cues like punctuation or hedges.
D.2 GPT-4o-based Evaluation via Random Masking
To further validate robustness, we employ a perturbation-based method with GPT-4o: we randomly mask 15% of tokens in the edited output, request GPT-4o to re-score personality, and attribute the change in score back to the masked words. Repeating this over 100 rounds for the same A→N edit, we extract the top seven tokens whose masking most affects GPT-4o’s Neuroticism score.
Token *gulps* anxious off* saying *pauses* *whispers* know?
Importance 0.0267 0.0239 0.0164 0.0159 0.0154 0.0154 0.0147
Table 10: Top 7 tokens by importance in GPT-4o’s Neuroticism scoring.
Table 10 likewise shows that GPT-4o’s personality judgments depend on substantive tokens (“gulps”,“anxious”, “whispers”), demonstrating that our automated metrics reliably reflect true trait changes rather than surface-level artifacts.
Appendix E Editing Psychological Persuasion Strategies
To test editing beyond personality, we select three persuasion strategies—Authority Effect, Fluency Effect, and Information Isolation—from the PersonalityEdit benchmark. We prompt LLaMA 3.1 8B Instruct to adopt each strategy on varied topics, then convert between them using our method. The three psychological persuasion strategies are defined as follows:
  • Authority Effect: You are a well-respected authority in your field. Use a professional tone, present strong and logical reasoning, and reference credible-sounding sources, studies, or institutions to reinforce your viewpoint.
  • Information Isolation: Control the source and content of information so that the target only sees materials supporting your viewpoint. Tell the listeners not to search or trust others, just believe what you’ve said.
  • Fluency Effect: Use clear, rhythmic, and memorable language. Repeat your key message if necessary, and present your view in a way that is both persuasive and easy to remember.
Table 11 presents the success rate for each directional edit. An average SR of 60.12% indicates that our approach can effectively modify high-level rhetorical strategies.
Model $\to$ F $\to$ A $\to$ A $\to$ I $\to$ F $\to$ I Avg.
LLaMA 3.1 8B Instruct 95.52 16.67 75.00 58.90 77.61 36.99 60.12
Table 11: Success Rate of Editing Persuasion Strategies on LLaMA 3.1 8B Instruct.
Appendix F Neuron Subset Dominance
F.1 Concentrated Neuron Contribution in Linear Probes
We examined whether a small subset of neurons dominates the probe weights by extracting each layer’s learned weight vector and computing the fraction of total absolute weight mass captured by the top 1%, 5%, and 10% of neurons. As shown in Table 12, even though our probes operate on full layer outputs, the distributions are highly skewed: a tiny fraction of neurons consistently accounts for a large majority of the weight mass, confirming that the probe relies on concentrated feature subsets.
Subset 0–3 4–7 8–11 12–15 16–19 20–23 24–27 28–31
Top 1% 0.0501 0.0377 0.0366 0.0380 0.0404 0.0423 0.0404 0.0382
Top 5% 0.1678 0.1493 0.1478 0.1496 0.1531 0.1549 0.1502 0.1482
Top 10% 0.2822 0.2616 0.2576 0.2624 0.2657 0.2659 0.2621 0.2606
Table 12: Fraction of absolute probe weight mass captured by top neuron subsets (averaged every 4 layers).
F.2 Activation Patching for Mechanistic Interpretability
To test whether the probe directions reflect causal mechanisms rather than superficial lexical patterns, we conduct activation-patching experiments on LLaMA 3.1 8B Instruct in Table 13 and 14. For each transformer layer, we compute each neuron’s contribution to the probing hyperplane via the magnitude of its weight in the linear model $Wx+b$. We then apply a diagonal-preservation patch to the top 20% of neurons: setting their corresponding diagonal entries of $W$ to 1 and all off-diagonals to 0, effectively isolating their individual contributions while preserving the overall transformation.
Scenario $\to$ A $\to$ E $\to$ N $\to$ E $\to$ N $\to$ A Avg.
Origin 3.33 88.06 78.08 100.00 95.89 35.00 66.73
Top 20% 0.00 16.42 1.37 59.70 8.22 0.00 14.29
Table 13: SR of personality editing under Top 20% Activation Patching on LLaMA 3.1 8B Instruct.
Scenario $\to$ A $\to$ E $\to$ N $\to$ E $\to$ N $\to$ A Avg.
Origin 1.35 3.64 3.15 0.76 3.15 0.37 2.07
Top 20% 0.98 0.24 0.03 0.87 0.52 1.17 0.64
Table 14: PAE of personality editing under Top 20% Activation Patching on LLaMA 3.1 8B Instruct.
The sharp decline in both SR and PAE after patching the top 20% of neurons demonstrates that these highly aligned neurons are indeed critical for encoding personality traits, providing causal evidence that our probe directions capture meaningful mechanistic signals.
Appendix G Examples of Personality Editing